# LQ-LoRA: Low-rank Plus Quantized Matrix Decomposition for Efficient Language Model Finetuning

### v2

@@ 3.2 Mixed-Configuration Quantization via an Integer Linear Program
Algorithm 1 $\operatorname{LQ-LoRA}$ (Section 3) $\{\mathbf{\mathbf{W}}^{(i)}\}_{i\in[N]}$: Parameters    $\{\mathbf{\mathbf{F}}^{(i)}\}_{i\in[N]}$:   Fisher information (optional)    $\mathcal{C}$: List of quantization configurations    $r$: LoRA rank    $B_{Q}$: Quantization budget # get quantization configurations (Section 3.2). $\{c^{(i)}\}\leftarrow\operatorname{GetConfig}(\{\mathbf{\mathbf{W}}^{(i)},\mathbf{\mathbf{F}}^{(i)}\},\mathcal{C},r,B_{Q})$ for $i\leftarrow 1$ to $N$ do      # matrix decomposition (Section 3.1).      $\mathbf{Q}^{(i)},\mathbf{L}_{1}^{(i)},\mathbf{L}_{2}^{(i)},\epsilon\leftarrow\operatorname{LQ}(\mathbf{W}^{(i)},\mathbf{F}^{(i)},c^{(i)},r)$$\{\mathbf{Q}^{(i)},\mathbf{L}_{1}^{(i)},\mathbf{L}_{2}^{(i)}\}_{i\in[N]}$ Algorithm 2 $\operatorname{LQ}$ (Section 3.1) $\mathbf{\mathbf{W}}$: Input weight matrix    $\mathbf{F}$: Fisher information (optional)    $c$: Quantization configuration    $r$: Target rank Initialize $\text{Initialize }\mathbf{Q}\leftarrow\mathbf{0}$ and $\epsilon_{0}\leftarrow\infty$ for $t\leftarrow 1$ to $T$ do      $\mathbf{L}_{1},\mathbf{L}_{2}\leftarrow\operatorname{Factorize}(\mathbf{W}-\mathbf{Q},\mathbf{F},r)$      $\mathbf{Q}\leftarrow\operatorname{Quantize}\left({\mathbf{W}-\mathbf{L}_{1}\mathbf{L}_{2}},c\right)$      if $\mathbf{F}$ is $\operatorname{None}$ then          $\epsilon_{t}\leftarrow\left\|\mathbf{W}{-}\left(\mathbf{Q}{+}\mathbf{L}_{1}\mathbf{L}_{2}\right)\right\|_{F}$      else          # weighted error (Section 3.3).          $\epsilon_{t}\leftarrow\left\|\sqrt{\mathbf{F}}\odot\left(\mathbf{W}{-}\left(\mathbf{Q}{+}\mathbf{L}_{1}\mathbf{L}_{2}\right)\right)\right\|_{F}$            if $\epsilon_{t}>\epsilon_{t-1}$ then break $\mathbf{Q},\mathbf{L}_{1},\mathbf{L}_{2},\epsilon_{t}$
Algorithm 3 $\operatorname{GetConfig}$ (Section 3.2) $\{\mathbf{\mathbf{W}}^{(i)}\}_{i\in[N]}$: Parameters    $\{\mathbf{\mathbf{F}}^{(i)}\}_{i\in[N]}$:   Fisher information (optional)    $\mathcal{C}$: List of quantization configurations    $r$: Target rank    $B_{Q}$: Quantization budget $\mathbf{E},\mathbf{S}\leftarrow\operatorname{zeros}(N,|\mathcal{C}|)$ initialize error and storage for $i\leftarrow 1$ to $N$ do      for $c\in\mathcal{C}$ do          $\mathbf{Q}^{(i)},\mathbf{L}_{1}^{(i)},\mathbf{L}_{2}^{(i)},\epsilon\leftarrow\operatorname{LQ}(\mathbf{W}^{(i)},\mathbf{F}^{(i)},c,r)$          $\mathbf{E}[i,c]\leftarrow\epsilon^{2}$          $\mathbf{S}[i,c]\leftarrow\operatorname{storage}(\mathbf{W}^{(i)},c)$       # get optimal configuration given budget with ILP. $\{c^{(i)}\}_{i\in[N]}\leftarrow\operatorname{ILPSolve}(\mathbf{S},\mathbf{E},B_{Q})$$\{c^{(i)}\}_{i\in[N]}$ Algorithm 4 $\operatorname{Factorize}$ (Section 3.3) $\mathbf{\mathbf{A}}$: Input matrix    $\mathbf{F}$: SVD weighting matrix (optional)    $r$: Target rank if $\mathbf{F}$ is $\operatorname{None}$ then      # (randomized) SVD with target rank $r$      $[\mathbf{U},\mathbf{\Sigma},\mathbf{V}^{\top}]\leftarrow\operatorname{SVD}(\mathbf{A},r)$      $\mathbf{L}_{1}\leftarrow\mathbf{U}\sqrt{\mathbf{\Sigma}}$, $\,\,\mathbf{L}_{2}\leftarrow\sqrt{\mathbf{\Sigma}}\mathbf{V}^{\top}$ else     # weighted SVD (Section 3.3).      $\mathbf{D}_{\text{row}}\leftarrow\operatorname{RowAverage}(\mathbf{F})$      $\mathbf{D}_{\text{col}}\,\leftarrow\operatorname{ColAverage}(\mathbf{F})$      $[\mathbf{U},\mathbf{\Sigma},\mathbf{V}^{\top}]\leftarrow\operatorname{SVD}(\mathbf{D}_{\text{row}}\mathbf{A}\mathbf{D}_{\text{col}},r)$      $\mathbf{L}_{1}\leftarrow\mathbf{D}_{\text{row}}^{-1}\mathbf{U}\sqrt{\mathbf{\Sigma}}$, $\,\,\mathbf{L}_{2}\leftarrow\sqrt{\mathbf{\Sigma}}\mathbf{V}^{\top}\mathbf{D}_{\text{col}}^{-1}$ $\mathbf{L}_{1},\mathbf{L}_{2}$
Dynamic quantization configurations.
@@ 4.2 LQ-LoRA for Model Compression
Table 3: LQ-LoRA comparison against other sub-4-bit PTQ methods. While we only experiment with LQ-LoRA on LLaMA-2 (bottom), we show other PTQ results on LLaMA-1 (top) as well to calibrate our results, as most prior works have focused on LLaMA-1. “Effective bits” takes into account the extra storage needed to store quantization parameters (e.g., scaling factors). In LQ-LoRA this includes the LoRA components, which are themselves quantized to 8 bits. For other methods, we take results corresponding to a setting with 3-bit quantization and a group-size 128 (if possible, otherwise the closest one). The effective bits for SpQR and LQ-LoRA are dependent on model size, and hence we show the effective bits for both settings. ${}^{\dagger}$esults from Shao et al. (2023). ${}^{\ddagger}$e show 3.1-bits instead of 3.01-bits with group size 128 because the latter performed worse.
@@
  RTN (3-bits, ${}^{\dagger}$ 3.15 8.62 6.10 7.01 4.24
  GPTQ (3-bits, g128(Frantar et al., ${}^{\dagger}$ 3.15 7.85 6.00 6.55 4.17
  AWQ (3-bits, g128(Lin et al., ${}^{\dagger}$ 3.15 7.92 5.94 6.46 3.99
@@
  OmniQuant (3-bits, g128(Shao et al., ${}^{\dagger}$ 3.15 7.75 5.93 6.15 3.94
  OmniQuant (2-bits, g64(Shao et al., ${}^{\dagger}$ 2.28 11.78 7.60 8.90 5.65
@@
LLaMA-2 ${}^{\dagger}$ 16 6.97 5.52 5.47 3.31
  RTN (3-bits, ${}^{\dagger}$ 3.15 8.40 6.02 6.66 3.97
  GPTQ (3-bits, g128(Frantar et al., ${}^{\dagger}$ 3.15 7.89 5.85 6.29 3.85
  AWQ (3-bits, g128(Lin et al., ${}^{\dagger}$ 3.15 7.84 - 6.24 -
  OmniQuant (3-bits, g128(Shao et al., ${}^{\dagger}$ 3.15 7.75 5.85 6.03 3.78
  OmniQuant (2-bits, g64(Shao et al., ${}^{\dagger}$ 2.28 12.72 7.88 9.62 6.11
@@ 5 Discussion and Limitations
We also discuss some negative results as well as limitations of LQ-LoRA. We found that re-factorizating the matrix periodically (e.g., after every $K$ gradient steps) did not yield improvements. Insofar as our initialization of $\mathbf{L}_{1}$ and $\mathbf{L}_{2}$ could orient the model to adapt itself in ways that may not be optimal for the task at hand, we also tried a hybrid approach where half of the adaptable low-rank component comes from LQ-LoRA, and the other half comes from standard LoRA initialization, but did not find this to improve results. Our approach also heavily relies on the fact that adaptation will occur through low-rank updates, and thus is not generally applicable to other parameter-efficient finetuning methods.
@@ 6 Related Work
Our work is closely related parameter-efficient finetuning. Popular methods include Adapters (Houlsby et al., 2019; Mahabadi et al., 2021), which insert trainable layers, prompt tuning (Li & Liang, 2021; Lester et al., 2021), which optimizes continuous prompts, and other methods which update subparts of the parameter vector (Guo et al., 2021; Zaken et al., 2022; Sung et al., 2021; Hu et al., 2022). (Some) parameter-efficient finetuning methods can reduce the GPU memory required for finetuning as there is no need to store the optimizer states associated with the fixed parameters. Recent work has combined parameter-efficient finetuning methods with quantization (Kwon et al., 2022; Dettmers et al., 2023a; Chai et al., 2023).
@@
Decomposing a data matrix into a low-rank matrix plus a sparse matrix (also known as robust PCA) is well-studied from both theoretical and applied perspectives (Lin et al., 2010; Zhou & Tao, 2011; 2013; Liu et al., 2013; Aravkin et al., 2014; Hintermuller & Wu, 2014; Yi et al., 2016; Zhang & Yang, 2017, inter alia). Within deep learning robust PCA has previously been applied to compress smaller models with fewer than 100M parameters (Chen & Ranftl, 2018; Cai et al., 2021). citetsaha2023matrix uses sketching techniques to obtain a quantized, low-rank approximation of a pretrained matrix. Recent contemporaneous work (Li et al., 2023) also performs low-rank plus quantized decomposition for LLM adaptation.
@@ Acknowledgements
We thank Minyoung Huh, Hongyi Wang, Isha Puri, and the members of the Hyundai 42dot research team for helpful comments and discussions. We are also grateful to Mengzhao Chen for clarification questions regarding OmniQuant. Eric Xing and Han Guo acknowledge the support of Microsoft PhD Fellowship, NGA HM04762010002, NSF IIS1955532, NSF CNS2008248, NIGMS R01GM140467, NSF IIS2123952, NSF BCS2040381, an Amazon Research Award, NSF IIS2311990, and DARPA ECOLE HR00112390063. This study was supported by funds from Hyundai Motor Group, MIT-IBM Watson AI, and the MLA@CSAIL initiative.
@@ Appendix A Implementation Details
Figure 5 displays the run-times of matrix multiplications between FP32 input data and quantized matrices. We dequantize the matrix just-in-time before executing the matmul, and hence the runtime is lower-bounded by FP32 matmul (dense). Notably, the dequantization overhead is relatively small for reasonably-sized matrices.
@@ Appendix B Full Results
Table 6 shows the full numeric results of LQ-LoRA vs. QLoRA and GPT-LoRA.

### v3

@@ 3.2 Mixed-Configuration Quantization via an Integer Linear Program
Algorithm 1 $\operatorname{LQ-LoRA}$ (Section 3)
$\{\mathbf{\mathbf{W}}^{(i)}\}_{i\in[N]}$: Parameters
   $\{\mathbf{\mathbf{F}}^{(i)}\}_{i\in[N]}$:   Fisher information (optional)
   $\mathcal{C}$: List of quantization configurations
   $r$: LoRA rank
   $B_{Q}$: Quantization budget
# get quantization configurations (Section 3.2).
$\{c^{(i)}\}\leftarrow\operatorname{GetConfig}(\{\mathbf{\mathbf{W}}^{(i)},\mathbf{\mathbf{F}}^{(i)}\},\mathcal{C},r,B_{Q})$
for $i\leftarrow 1$ to $N$ do
     # matrix decomposition (Section 3.1).
     $\mathbf{Q}^{(i)},\mathbf{L}_{1}^{(i)},\mathbf{L}_{2}^{(i)},\epsilon\leftarrow\operatorname{LQ}(\mathbf{W}^{(i)},\mathbf{F}^{(i)},c^{(i)},r)$
$\{\mathbf{Q}^{(i)},\mathbf{L}_{1}^{(i)},\mathbf{L}_{2}^{(i)}\}_{i\in[N]}$
Algorithm 2 $\operatorname{LQ}$ (Section 3.1)
$\mathbf{\mathbf{W}}$: Input weight matrix
   $\mathbf{F}$: Fisher information (optional)
   $c$: Quantization configuration
   $r$: Target rank
Initialize $\text{Initialize }\mathbf{Q}\leftarrow\mathbf{0}$ and $\epsilon_{0}\leftarrow\infty$
for $t\leftarrow 1$ to $T$ do
     $\mathbf{L}_{1},\mathbf{L}_{2}\leftarrow\operatorname{Factorize}(\mathbf{W}-\mathbf{Q},\mathbf{F},r)$
     $\mathbf{Q}\leftarrow\operatorname{Quantize}\left({\mathbf{W}-\mathbf{L}_{1}\mathbf{L}_{2}},c\right)$
     if $\mathbf{F}$ is $\operatorname{None}$ then
         $\epsilon_{t}\leftarrow\left\|\mathbf{W}{-}\left(\mathbf{Q}{+}\mathbf{L}_{1}\mathbf{L}_{2}\right)\right\|_{F}$
     else
         # weighted error (Section 3.3).
         $\epsilon_{t}\leftarrow\left\|\sqrt{\mathbf{F}}\odot\left(\mathbf{W}{-}\left(\mathbf{Q}{+}\mathbf{L}_{1}\mathbf{L}_{2}\right)\right)\right\|_{F}$
     if $\epsilon_{t}>\epsilon_{t-1}$ then break
$\mathbf{Q},\mathbf{L}_{1},\mathbf{L}_{2},\epsilon_{t}$
Algorithm 3 $\operatorname{GetConfig}$ (Section 3.2)
$\{\mathbf{\mathbf{W}}^{(i)}\}_{i\in[N]}$: Parameters
   $\{\mathbf{\mathbf{F}}^{(i)}\}_{i\in[N]}$:   Fisher information (optional)
   $\mathcal{C}$: List of quantization configurations
   $r$: Target rank
   $B_{Q}$: Quantization budget
$\mathbf{E},\mathbf{S}\leftarrow\operatorname{zeros}(N,|\mathcal{C}|)$ initialize error and storage
for $i\leftarrow 1$ to $N$ do
     for $c\in\mathcal{C}$ do
         $\mathbf{Q}^{(i)},\mathbf{L}_{1}^{(i)},\mathbf{L}_{2}^{(i)},\epsilon\leftarrow\operatorname{LQ}(\mathbf{W}^{(i)},\mathbf{F}^{(i)},c,r)$
         $\mathbf{E}[i,c]\leftarrow\epsilon^{2}$
         $\mathbf{S}[i,c]\leftarrow\operatorname{storage}(\mathbf{W}^{(i)},c)$
# get optimal configuration given budget with ILP.
$\{c^{(i)}\}_{i\in[N]}\leftarrow\operatorname{ILPSolve}(\mathbf{S},\mathbf{E},B_{Q})$
$\{c^{(i)}\}_{i\in[N]}$
Algorithm 4 $\operatorname{Factorize}$ (Section 3.3)
$\mathbf{\mathbf{A}}$: Input matrix
   $\mathbf{F}$: SVD weighting matrix (optional)
   $r$: Target rank
if $\mathbf{F}$ is $\operatorname{None}$ then
     # (randomized) SVD with target rank $r$
     $[\mathbf{U},\mathbf{\Sigma},\mathbf{V}^{\top}]\leftarrow\operatorname{SVD}(\mathbf{A},r)$
     $\mathbf{L}_{1}\leftarrow\mathbf{U}\sqrt{\mathbf{\Sigma}}$, $\,\,\mathbf{L}_{2}\leftarrow\sqrt{\mathbf{\Sigma}}\mathbf{V}^{\top}$
else     # weighted SVD (Section 3.3).
     $\mathbf{D}_{\text{row}}\leftarrow\operatorname{RowAverage}(\mathbf{F})$
     $\mathbf{D}_{\text{col}}\,\leftarrow\operatorname{ColAverage}(\mathbf{F})$
     $[\mathbf{U},\mathbf{\Sigma},\mathbf{V}^{\top}]\leftarrow\operatorname{SVD}(\mathbf{D}_{\text{row}}\mathbf{A}\mathbf{D}_{\text{col}},r)$
     $\mathbf{L}_{1}\leftarrow\mathbf{D}_{\text{row}}^{-1}\mathbf{U}\sqrt{\mathbf{\Sigma}}$, $\,\,\mathbf{L}_{2}\leftarrow\sqrt{\mathbf{\Sigma}}\mathbf{V}^{\top}\mathbf{D}_{\text{col}}^{-1}$
$\mathbf{L}_{1},\mathbf{L}_{2}$
@@ 4.2 LQ-LoRA for Model Compression
Table 3: LQ-LoRA comparison against other sub-4-bit PTQ methods. While we only experiment with LQ-LoRA on LLaMA-2 (bottom), we show other PTQ results on LLaMA-1 (top) as well to calibrate our results, as most prior works have focused on LLaMA-1. “Effective bits” takes into account the extra storage needed to store quantization parameters (e.g., scaling factors). In LQ-LoRA this includes the LoRA components, which are themselves quantized to 8 bits. For other methods, we take results corresponding to a setting with 3-bit quantization and a group-size 128 (if possible, otherwise the closest one). The effective bits for SpQR and LQ-LoRA are dependent on model size, and hence we show the effective bits for both settings. Results from Shao et al. (2023). We show 3.1-bits instead of 3.01-bits with group size 128 because the latter performed worse.
@@
RTN (3-bits, g128) 3.15 8.62 6.10 7.01 4.24
GPTQ (3-bits, g128(Frantar et al., 2022) 3.15 7.85 6.00 6.55 4.17
AWQ (3-bits, g128(Lin et al., 2023) 3.15 7.92 5.94 6.46 3.99
@@
OmniQuant (3-bits, g128(Shao et al., 2023) 3.15 7.75 5.93 6.15 3.94
OmniQuant (2-bits, g64(Shao et al., 2023) 2.28 11.78 7.60 8.90 5.65
@@
LLaMA-2 Uncompressed 16 6.97 5.52 5.47 3.31
RTN (3-bits, g128) 3.15 8.40 6.02 6.66 3.97
GPTQ (3-bits, g128(Frantar et al., 2022) 3.15 7.89 5.85 6.29 3.85
AWQ (3-bits, g128(Lin et al., 2023) 3.15 7.84 - 6.24 -
OmniQuant (3-bits, g128(Shao et al., 2023) 3.15 7.75 5.85 6.03 3.78
OmniQuant (2-bits, g64(Shao et al., 2023) 2.28 12.72 7.88 9.62 6.11
@@ 5 Discussion and Limitations
We also discuss some negative results as well as limitations of LQ-LoRA. We found that re-factorizing the matrix periodically (e.g., after every $K$ gradient steps) did not yield improvements. Insofar as our initialization of $\mathbf{L}_{1}$ and $\mathbf{L}_{2}$ could orient the model to adapt itself in ways that may not be optimal for the task at hand, we also tried a hybrid approach where half of the adaptable low-rank component comes from LQ-LoRA, and the other half comes from standard LoRA initialization, but did not find this to improve results. Our approach also heavily relies on the fact that adaptation will occur through low-rank updates, and thus is not generally applicable to other parameter-efficient finetuning methods.
@@ 6 Related Work
Our work is closely related to parameter-efficient finetuning. Popular methods include Adapters (Houlsby et al., 2019; Mahabadi et al., 2021), which insert trainable layers, prompt tuning (Li & Liang, 2021; Lester et al., 2021), which optimizes continuous prompts, and other methods which update subparts of the parameter vector (Guo et al., 2021; Zaken et al., 2022; Sung et al., 2021; Hu et al., 2022). (Some) parameter-efficient finetuning methods can reduce the GPU memory required for finetuning as there is no need to store the optimizer states associated with the fixed parameters. Recent work also combines parameter-efficient finetuning methods with quantization (Kwon et al., 2022; Dettmers et al., 2023a).
@@
Decomposing a data matrix into a low-rank matrix plus a sparse matrix (also known as robust PCA) is well-studied from both theoretical and applied perspectives (Lin et al., 2010; Zhou & Tao, 2011; 2013; Liu et al., 2013; Aravkin et al., 2014; Hintermuller & Wu, 2014; Yi et al., 2016; Zhang & Yang, 2017, inter alia). Within deep learning robust PCA has previously been applied to compress smaller models with fewer than 100M parameters (Chen & Ranftl, 2018; Cai et al., 2021). Saha et al. (2023) uses sketching techniques to obtain a quantized, low-rank approximation of a pretrained matrix. Recent contemporaneous work (Li et al., 2023) also performs low-rank plus quantized decomposition for LLM adaptation.
@@ Acknowledgements
We thank Minyoung Huh, Isha Puri, Hongyi Wang, Lirui Wang, and the members of the Hyundai 42dot research team for helpful comments and discussions. We are also grateful to Mengzhao Chen for clarification questions regarding OmniQuant. Eric Xing and Han Guo acknowledge the support of Microsoft PhD Fellowship, NGA HM04762010002, NSF IIS1955532, NIGMS R01GM140467, NSF IIS2123952, NSF BCS2040381, NSF IIS2311990, SRC AIHW 2024AH3210, and DARPA ECOLE HR00112390063. This study was supported by funds from Hyundai Motor Group, MIT-IBM Watson AI, and the MLA@CSAIL initiative.
@@ Appendix A Implementation Details
Figure 5 displays the run-times of matrix multiplications between FP32 input data and quantized matrices. We dequantize the matrix just-in-time before executing the matmul, and hence the runtime is lower-bounded by FP32 matmul (dense). We enable TF32 in PyTorch to utilize Tensor Cores, and collect CUDA time through PyTorch Profiler. Notably, the dequantization overhead is relatively small for reasonably-sized matrices.
@@ Appendix B Full Results
Table 6 shows the full numeric results of LQ-LoRA vs. QLoRA and GPTQ-LoRA.